OpenStudy (anonymous):

Air is being pumped into a spherical balloon at a rate of 20ft 3/min. At what rate is the radius changing when the radius is 3 ft?

OpenStudy (amistre64):

it asks to find dr/dt it gives as dV/dt = 20; so lets equate volume with radius V = (4/3) pi r^3 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

when we derive both sides we get: \[\frac{dV}{dt} = 4pi\ r^2 \frac{dr}{dt}\] we need to solve for dr/dt

OpenStudy (amistre64):

20 = 4pi (3)^2 * dr/dt 20/(36pi) = dr/dt right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

got it: .176

OpenStudy (anonymous):

right?

OpenStudy (amistre64):

Another way to look at these is thru the chain rule: \[\frac{dr}{dt}=\frac{dr}{dV}\frac{dV}{dt}\] and then you can focus on getting dr/dV V = (4pi/3) r^3 ; derive with respect to V dV/dV = (4pi) r^2 dr/dV ; dV/dV = 1 1/(4pi r^2) = dr/dV ; at r=3 we get 1/(36pi) = dr/dV \[\frac{dr}{dt}=\frac{1}{36pi}\frac{20}{1}\]

OpenStudy (amistre64):

.1768... yes